Ncert Solutions Maths class 12th

Get insights from 2.5k questions on Ncert Solutions Maths class 12th, answered by students, alumni, and experts. You may also ask and answer any question you like about Ncert Solutions Maths class 12th

  • 2.5k

    Questions

  • 0

    Discussions

  • 1

    Active Users

  • 65

    Followers

Have a question about Ncert Solutions Maths class 12th? Ask our experts

Ask Your Question

Keep it short & simple. Type complete word. Avoid abusive language.

New answer posted•a year ago
S
a year ago

34. P (A′B′)= [1−P (A)] [1−P (B)]

⇒P (A′∩B′)=1−P (A)−P (B)+P (A)P (B)

⇒− [P (A)+P (B)−P (A∩B)]=−P (A)−P (B)+P (A) (B)

⇒−P (A)−P (B)+P (A∩B)=−P (A)−P (B)+P (A)P (B)

⇒P (A∩B)=P (A).P (B)

Therefore, it shows A and B are independant

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

33. The outcome of rolling 2 dice is 36 . The only prime number is 2 .

Let,  E= event of getting an even prime number on each die

E= { (2, 2)} therefore

P (E)=136

Therefore, the correct answer is  (D) 136

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

32. Let,

P(H)= denote the student who read Hindi newspaper

P(E)= denote the student who read English newspaper

Then, P(H)=60%=610=35

P(E)=40%=410=25

i. P(H∩E)=20%=210=15

⇒P(H∪E)′=1−[P(H)+P(E)−P(H∩E)]

=1−[35+25−15]

=1−3+2−15

=1−45=15

ii. Probability of randomly chosen student that reads English newspaper, if she reads Hindi newspaper, is given by

P(E/H)=P(E∩H)P(H)=1535=13

iii. Probability that randomly chosen student reads Hindi newspaper

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

31. i. In a deck of 52 cards,

13 are spades and 4 are aces.

∴P(E)=P (the card drawn is spade) =1352=14

∴P(F)=P (the card drawn is an ace) =452=113

∴P(E∩F)=P (the card drawn is spade and an ace) =152

P(E).P(F)=14×113=152=P(E∩F)

⇒P(E).P(F)=P(E∩F)


E and F are independent.

ii. In a deck of 52 cards,

26 cards are black and 4 a

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

30. Let, P(A)= probability of solving problem by A=12

P(B)= probability of solving problem by B=13

Since, the problem is solved independently by A and B , then,

⇒P(A∩B)=P(A).P(B)

⇒P(A∩B)=12×13

=16

P(A′)=1−P(A)=1−12=12

P(B′)=1−P(B)=1−13=23

i. Probability of the problem is solved

⇒P(A∪B)=P(A)+P(B)−P(A∩B)

⇒P(A∪B)=12+13−16

=3+2−16

=46=23

ii. Probability that exactly one of the solved problem is given by P(A)P(B′)+P(B)P(A′)=12×23+12×13

=26+16=36=12

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

29. Total number of balls =18

Number of red balls =8

Number of black balls =10

i. Probability of getting red ball at first draw =818=49

The red ball is replaced after the first draw

∴ Probability of getting red ball in second draw =818=49

∴ Probability of getting both the balls red =49×49=1681

ii. Probability o

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

28. The sample space of given condition is

S= {1, 2, 3, 4, 5, 6}

Let,  P (A)= probability of getting an odd number in first throw

⇒P (A)=36=12

P (B)= probability of getting an even number

⇒P (B)=36=12

Probability of getting an even number in three times =12×12×12=18

So, Probability of getting an odd number at least once

=1− probability of getting an odd

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

27. Given,

P(A)=0.3

P(B)=0.6

i. A and B are independent

⇒P(A∩B)=P(A).P(B)

=0.3×0.6

=0.18

ii. P ( A and not B )

⇒P(A∩B′)=P(A)−P(A∩B)

=0.3−0.18

=0.12

iii. P ( A or B )

⇒P(A∪B)=P(A)+P(B)−P(A∩B)

=0.3+0.6−0.18

=0.72

iv. P (neither A nor B )

⇒P(A′∩B′)=P(A∪B)′

=1−P(A∪B)

=1−0.72

=0.28

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

26. Given,

P(A)=12

P(B)=712

P (not A or not B )= 14

⇒P(A′∪B′)=P(A∩B)′

⇒P(A∩B)′=14

⇒1−P(A∩B)=14

⇒P(A∩B)=1−14

⇒4−14

⇒34

Now,

P(A).P(B)=12×712

=724≠34

⇒P(A).P(B)≠P(A∩B)

Therefore, A and B are not independent.

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

25. Given,

P (A)=14

P (B)=12

P (A∩B)=18 Here,  P  (not A not B ) =P (A′∩B′)=P (A∪B)′

⇒1−P (A∪B)

⇒1− [P (A)+P (B)−P (A∩B)]

⇒1− [14+12−18]

⇒1− [2+4−18]

⇒1−58=8−58

⇒38

Thumbs Up IconUpvote Thumbs Down Icon